The price of a commodity increased from rs40 to rs6. The quantity demanded decreased from 25 units to 15 units. Calculate arc elasticity of demand and own price elasticity of demand. What is the degree of elasticity of demand.


1
Expert's answer
2021-02-11T13:02:00-0500

"Arc, E_d=[\frac{(Q_2-Q_1)}{midpoint Q_d}]\div[\frac{(P_2-P_1)}{midpoint P}]"

"=[\frac{(15-25)}{(25+15)}1/2]\div[\frac{(60-40)}{(60+40)}1/2]=\frac{-10}{20}\div \frac{20}{50}=-0.5\div 0.4"

"=|-1.25|=1.25"

own price of demand"=\%" change in "Q_d" "/ \%" change in "P=\frac{(15-25)}{25} \times100\%\div \frac{(60-40)}{40}\times100\%=-40\div50=|-0.8|=0.8"

degree of elasticity "=\frac{\Delta Q_d}{Delta P}=\frac{(15-25)}{(60-40)}=\frac{-10}{20}=|-0.5|=0.5"